150,000,000 km in scientific notation?
3961 km in scientific notation?
1.63 m in scientific notation?

Answer by CubeyThePenguin(3113)

150,000,000 km = 1.5 * 10^(8) km
3961 km = 3.961 * 10^(3) km
1.63 m = 1.63 * 10^(0) km